# Helicity controls the direction of fluxes in rotating turbulence

### v1

@@ Abstract
In sector H, species $s$ and s-s are uncoupled and $\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}=0$. Equation 5 therefore becomes a continuity equation for $H_{\bm{p}}^{s}$: helicity in sector H is conserved by sign, which is equivalent to the conservation of wave action $\langle|a_{\bm{p}}^{s}|^{2}\rangle/\omega_{\bm{p}}^{s}$ for each species $s$. Due to this additional sign-definite invariant, the steady-state energy flux solving (5) decays with $p_{y}$: $\Pi^{s}_{\bm{p}}=\epsilon^{s}_{p_{x},p_{z}}k_{f}/p$ for $p>k_{f}$. As $\epsilon^{s}_{p_{x},p_{z}}-\Pi^{s}_{|p_{y}|\to\infty}>0$, this implies that energy is transferred to the large-scale 2D flow, as depicted in Fig. 2.
At large-enough rotation, all modes are effectively in sector $H$. In this regime, all waves interacting with the large-scale 2D flow conserve their single-sign helicity: rotation has turned sign-definite an otherwise sign-indefinite invariant. This regime is obtained for $Ro_{\epsilon}\lesssim 0.1$ in Fig. 1(d), for which the transfer of energy is purely inverse, going from the waves to the large-scale 2D flow [gome2025waves, kolokolov2020structure]. In contrast, for $Ro_{\epsilon}\gtrsim 0.1$, heterochiral interactions are possible and generate a finite flux of energy to small scales, as we now explain.
Figure 2: (a) Energy transfer to the large-scale 2D flow in spectral space $(p_{y},p_{z})$ from QL theory and (b) schematic of the energy fluxes in the system. Orange line in (a) marks the boundary between Sectors H and A, (4). Waves in sector H conserve their sign-definite helicity $H_{\bm{p}}^{\pm}$, constraining the forward energy flux and causing energy transfer to the 2D flow. In contrast, modes in sector A break this conservation and extract energy from the 2D flow, with resulting energy flux $\Pi>\Pi^{*}$ (6), $\Pi^{*}$ being the energy input from sector H.
Energy flux in sector A — Modes in sector A are shear-dominated, and their energy balance is typical for a non-rotating flow. There, modes of opposite helicities interact, helicity is no longer conserved by sign, and energy is extracted from the shear flow. Using scale separation $l_{f}/L_{y}\ll 1$, the dynamics effectively reduces to that for a linear shear flow, which is analytically solvable when changing variables to fluctuating cross-flow velocity $v$ and vorticity $\eta$, see Eq. (12) and [farrell1993optimal]. The condensate generates vorticity from $z$-varying $v$, which tilts the mean shear (via the RHS of Eq. (12b)), and converts mean-flow energy into turbulent kinetic energy (via Eq. (12c)).
Assuming an incoming flux of $\Pi^{*}(p_{x},p_{z})$ at $p=p^{*}$ into sector A, the flux $\Pi_{\bm{p}}\equiv\Pi^{+}_{\bm{p}}+\Pi^{-}_{\bm{p}}$ can be exactly computed (see End Matter), reading
$\Pi_{\bm{p}}$ $\underset{p_{y}\gg p^{*}_{y}}{=}\frac{\Pi^{*}}{2}\Bigg[1+\frac{p^{*2}}{p_{0}^{2}}\frac{p_{z}^{2}}{p_{x}^{2}}\Big(\frac{\pi}{2}-\arctan\Big(\frac{p_{y}^{*}}{p_{0}}\Big)\Big)^{2}\Bigg],$ (6)
@@
Total energy transfer and closure — To combine the results for sector H and A we, take the incoming flux $\Pi^{*}=\epsilon_{p_{x},p_{z}}(p_{x},p_{z})k_{f}/p^{*}$. The energy transfer to the 2D condensate as a function of the wavenumber $\bm{p}$ of inertial waves, ${U^{\prime}}\langle uv\rangle_{\bm{p}\bm{k}}$, can then be computed for a fixed value of ${U^{\prime}}/\Omega$ (see Eq. (19) and Fig. 2(a)). Waves in sector $H$ transfer energy to the condensate (${U^{\prime}}\langle uv\rangle_{\bm{p}\bm{k}}>0$), producing a decaying energy flux, which flows into sector A, where waves extract energy from the condensate (${U^{\prime}}\langle uv\rangle_{\bm{p}\bm{k}}<0$), see the schematic Fig. 2(b).
@@
Figure 3: (a) Energy transfer from 3D waves to 2D condensate and (b) rescaled amplitude of the condensate, as measured from the DNS and predicted from the QL theory. Dashed lines show the predictions in each regime: in black, Eq. (8); in red, the scalings derived in [gome2025waves]. Solid lines show a numerical solution of the QL system at $Re=46$, including a wave cutoff $k_{\Omega}$ and a smooth treatment of homochiral-wave near-resonances. Inset in (b): rescaled amplitude of the 2D flow as a function of $Re$, showing the saturation to a flux-loop state.
@@
For $Ro_{\epsilon}\ll 1/b\sim 8l_{f}/L_{x}$ (here $\simeq 1.2$ when including subleading corrections), Eq.(8) gives $T_{\rm 3D-2D}\simeq\epsilon$ and $U^{\prime}\simeq\sqrt{\epsilon/\nu}$, corresponding to almost purely homochiral-wave interactions. This is consistent with the DNS behavior in Fig. 1 for $Ro_{\epsilon}\lesssim 0.1$. For smaller $Ro_{\epsilon}$ the resonance condition (3) restricts homochiral-wave interactions, leading to a decoupling between 3D waves and the 2D flow, see the red dashed line in Fig. 3(a) obtained in [gome2025waves]. Note that, in the theory, the transition from the purely-homochiral to the mixed-interaction regime around $Ro_{\epsilon}\simeq 0.1$ is smoothed out when using a smooth function of ${U^{\prime}}/\Omega$ (21) instead of the sharp near-resonant condition (3) for $\tilde{s}=s$. This perfects the agreement with the DNS in this region (solid line in Fig. 3(a)).
At the other end, for $Ro_{\epsilon}=\text{Ro}\times\text{Re}^{1/2}\gg 1/b$, Eq. (8) predicts that $U^{\prime}/\Omega\to 1/b$, a Reynolds-independent value. This asymptotic state corresponds to a vanishing energy transfer $T_{\rm 3D-2D}\to 0^{+}$. For fixed Ro and increasing Re in the DNS, ${U^{\prime}}/\Omega$ is indeed found to saturate to a constant value, closely following this prediction for the lowest $Ro$ (inset of Fig. 3b), and in accordance with previous observations of rotating turbulence under 3D [alexakis2015rotating] or 2D forcing [seshasayanan2018condensates].
The asymptotic flow state $T_{\rm 3D-2D}\underset{Re\to\infty}{\to}0^{+}$ with finite $U^{\prime}$ is dubbed a flux-loop state: a state where inverse and forward energy transfers are almost counterbalanced, see also [boffetta2011flux, falkovich2017vortices, seshasayanan2018condensates, de2024pattern, clark2020phase]. This behavior, consistent with taking $\nu\to 0$ in the energy balance (1), would be impossible if the system strictly conserved two sign-definite quantities, as no steady state would be reached for $Re\to\infty$. Here, we reveal that the condensate self-tunes to a state of zero large-scale dissipation in this limit, achieved by an exact balance between the inverse energy transfer from modes conserving $H_{\bm{p}}^{\pm}$ (sector H), and the forward transfer due to modes breaking this conservation (sector A). Note that viscosity (and so Re) here primarily affects the large-scale dissipation, and could be replaced by, e.g., Ekman friction [le2017inertial].
@@
When, instead, fixing Re and increasing Ro, scaling the $U^{\prime}\sim\Omega$ implies a vanishing 2D flow only for zero rotation $(Ro=\infty$), while condensates in DNS disappear to eddy turbulence at $Ro\simeq 0.5$, see Fig. 4. This is due to a shortage of waves at low rotation, and hence of modes in sector H, as we now explain.
When decreasing rotation, more modes behave as eddies instead of inertial waves, i.e. are such that $\tau_{nl}^{\rm eddy}<\tau_{\Omega}\equiv 2/\omega_{\bm{p}}$ (or, equivalently, $p>k_{\Omega}\equiv\Omega^{3/5}\epsilon^{-1/5}p_{z}^{3/5}$ [zeman1994note]), with $\tau_{nl}^{\rm eddy}=\epsilon^{-1/3}p^{-2/3}$ the Kolmogorov eddy turnover time. Modes belonging to sector H are such that $\tau_{\Omega}\lesssim 1/{U^{\prime}}$, following Eq. (4). For $p>k_{\Omega}$, where they behave like eddies and $\tau_{nl}^{\rm eddy}<\tau_{\Omega}$, modes do not interact dominantly with the condensate as $\tau_{nl}^{\rm eddy}<1/{U^{\prime}}$, which breaks the QL approximation. As a consequence, the homochiral sector H, which is the source of energy for the condensate, is cutoff at $p=k_{\Omega}$ (sector A being cutoff at $k_{U}$ where eddy-eddy interactions dominate). As rotation is decreased, $k_{\Omega}$ diminishes and the energy input to the condensate is reduced, until no modes exist in sector H when $Ro=0.5$, for which $k_{\Omega}(p_{z}=k_{f})=k_{f}$. This marks the disappearance of condensates to a strong turbulence of 3D eddies. Introducing the wavenumber cutoff into our QL theory and integrating the resulting Reynolds stress numerically, we reproduce the drop in energy transfer and condensate amplitude observed in DNS, see the solid lines in Fig. 3.
@@
While jets are found for most $Ro$ and $Re$, the condensate takes the form of cyclonic vortices at the boundary with 3D turbulence, (stars in Figs. 1 and 4), previously observed in [marchetti2025spontaneous, seshasayanan2018condensates, clark2020phase]. Our theory roughly captures the condensate amplitude for these states, but cannot explain scale selection nor symmetry breaking. The former may be due to time-scale competition between modes of different scales, consistent with the metastability of such states [estrada2025stability].
@@
We have untangled the organization of kinetic energy fluxes in rotating turbulence, showing it is controlled by helicity. Large-scale 2D structures are energized by fast-enough inertial waves via wave-dominated interactions which conserve helicity by sign. In contrast, the 2D flow loses energy to slow waves through shear-dominated interactions which break helicity-by-sign conservation, the waves carrying energy to small scales similarly to non-rotating 3D turbulence. We have determined these bi-directional energy transfers analytically across rotation regimes via a quasi-linear wave-kinetic theory, which reproduces the $Ro$ and $Re$ behavior of 2D condensates in Navier-Stokes simulations, as summarized in Fig. 4.
Our results show that large-scale self-organization can be regulated by a sign-indefinite invariant in the presence of waves [biferale2012inverse, shavit2024sign], and thus arise even in systems that do not conserve multiple sign-definite quantities. The mean-turbulent framework developed here provides a route to analyzing such behavior in other settings – such as thin layers [benavides2017critical], chiral flows [de2024pattern], stratified flows [chini2022exploiting, shah2024regimes, shavit2024sign, labarre20242d], magnetohydrodynamics [pouquet2019helicity, gallet2012reversals] plasmas [diamond2005zonal] or geophysical flows [guillon2025self] — and to elucidating the energy sinks and feedback mechanisms governing their self-organization.
@@ Acknowledgements.
We detail here the governing equations and solutions to the quasi-linear system describing the interactions between 3D inertial waves and a large-scale 2D mean flow. We consider a sinusoidal condensate of single mode $\bm{k}=\pm 2\pi/L_{y}$, $U_{\bm{k}}={U^{\prime}}/\sqrt{2}\delta_{\bm{k}-2\pi/L_{y}}$. The fluctuation field $\bm{u}(\bm{x})$ is decomposed into the helical basis $\bm{h}_{\bm{p}}^{s}=(\bm{p}\times[\bm{p}\times\mathbf{e}_{z}]-isp\bm{p}\times\mathbf{e}_{z})/(\sqrt{2}pp_{\perp})$, with $p=|\bm{p}|$ and $p_{\perp}=\sqrt{p_{x}^{2}+p_{y}^{2}}$. Following wave kinetic theory, we obtain evolution equations for the wave correlators $\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s^{\prime}}\rangle$, which we derive from the wave dynamical equation (2) in an expansion in ${U^{\prime}}/\Omega$ (see [gome2025waves] and SM §B), and result in the helicity balance (5).
In Sector H, waves conserve their single-sign helicity, via Eq. (5) with $\langle uv\rangle_{\bm{p}\bm{k}}^{s,-s}=0$. Introducing the correlator $\Phi_{\bm{p}\bm{k}}^{s}\equiv\frac{1}{2}(\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{s}\rangle)$, Eq. (5) is solved by the energy flux
$\Pi_{\bm{p}}^{s}\equiv\frac{-{U^{\prime}}p_{x}}{\sqrt{2}}\Phi_{\bm{p}\bm{k}}^{s}=\left\{\begin{aligned} &~0~~\text{ if }\sigma p_{y}<-p_{y}^{f}\\ &\frac{\sigma\epsilon^{s}_{p_{x},p_{z}}}{2}\frac{k_{f}}{p}~~\text{ if }-p_{y}^{f}<\sigma p_{y}<p_{y}^{f}\\ &\sigma\epsilon^{s}_{p_{x},p_{z}}\frac{k_{f}}{p}~~\text{ if }\sigma p_{y}>p_{y}^{f},\end{aligned}\right$. (9)
@@
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\bm{p}\bm{k}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\frac{p_{x}p_{y}}{p^{3}}(\Phi_{\bm{p}\bm{k}}^{+}+\Phi_{\bm{p}\bm{k}}^{-})$ (10)
which is strictly positive when $\sigma p_{y}>p_{y}^{f}$, and provided that the interaction is resonant at $p=k_{f}$: $\omega_{\bm{p}}^{s}+\omega_{-\bm{p}-\bm{k}}^{s}<{U^{\prime}}$.
In Sector A, the 2D flow mediates interactions between modes of opposite helicities, hence correlators $\Psi_{\bm{p}\bm{k}}^{s}\equiv\frac{1}{2}(\langle a_{\bm{p}}^{s}a_{-\bm{p}+\bm{k}}^{-s}\rangle+\langle a_{\bm{p}}^{s}a_{-\bm{p}-\bm{k}}^{-s}\rangle)$ are generated. They are determined from an evolution equation for $\langle a_{\bm{p}}^{s*}a_{\bm{p}}^{-s}\rangle$:
$\partial_{t}\langle a_{\bm{p}}^{s*}a_{\bm{p}}^{-s}\rangle=$ $\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\mathcal{L}_{\bm{p}}[\Psi_{\bm{p}\bm{k}}^{s,-s*}+\Psi_{\bm{p}\bm{k}}^{-s,s}])$
$~~~~~~-2V_{\bm{p}}^{s,-s*}(\Phi_{\bm{p}\bm{k}}^{s*}+\Phi_{\bm{p}\bm{k}}^{-s})\Bigg),$ (11)
where $\mathcal{L}_{\bm{p}}\equiv\Big[p_{x}\partial_{p_{y}}+\frac{p_{x}p_{y}}{p^{2}}-i\frac{p_{z}}{p}(1+2\frac{p_{x}^{2}}{p_{\perp}^{2}})\Big]$. Together with Eq. (5), Eq. (11) constitutes a closed system of correlators $(\Phi_{\bm{p}\bm{k}}^{s},\Psi_{\bm{p}\bm{k}}^{s})$. By changing variables from helical amplitudes $(a_{\bm{p}}^{+},a_{\bm{p}}^{-})$ to cross-flow velocity $v_{\bm{p}}$ and vorticity $\eta_{\bm{p}}\equiv(\nabla\times\bm{u})_{\bm{p},y}$, we obtain a triangular system of correlators from Eqs. (5)-(11), written in steady state as:
$\partial_{p_{y}}\Big(p^{4}\Re(\langle vv\rangle_{\bm{p}\bm{k}})\Big)=-p^{4}\frac{\sqrt{2}\epsilon\chi_{\bm{p},y}}{p_{x}{U^{\prime}}}$ (12a)
$\partial_{p_{y}}\Big(p^{2}\langle\eta v\rangle_{\bm{p}\bm{k}}\Big)=ip^{2}\frac{p_{z}}{p_{x}}\langle vv\rangle_{\bm{p}\bm{k}}$ (12b)
$\partial_{p_{y}}\Re(\langle\eta\eta\rangle_{\bm{p}\bm{k}})=\frac{2p_{z}}{p_{x}}\Im\langle\eta v\rangle_{\bm{p}\bm{k}}-\frac{\sqrt{2}\epsilon p^{2}\chi_{\bm{p},y}}{p_{x}{U^{\prime}}},$ (12c)
with $\langle fg\rangle_{\bm{p}\bm{k}}\equiv(\langle f_{\bm{p}}g_{-\bm{p}+\bm{k}}\rangle+\langle f_{\bm{p}}g_{-\bm{p}-\bm{k}}\rangle)/2$. As the 3D modes have low frequencies in Sector A, rotation has no effect on the leading-order dynamics. Therefore, system (12) describes a non-rotating shear flow interacting with turbulent fluctuations.

### v2

@@ Abstract
In sector H, species $s$ and s-s are uncoupled and $\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}=0$. Equation 5 therefore becomes a continuity equation for $H_{\bm{p}}^{s}$: helicity in sector H is conserved by sign, which is equivalent to the conservation of wave action $\langle|a_{\boldsymbol{p}}^{s}|^{2}\rangle/\omega_{\boldsymbol{p}}^{s}$ for each species $s$. Due to this additional sign-definite invariant, the steady-state energy flux solving (5) decays with $p_{y}$: $\Pi^{s}_{\boldsymbol{p}}=\epsilon^{s}_{p_{x},p_{z}}k_{f}/p$ for $p>k_{f}$. As $\epsilon^{s}_{p_{x},p_{z}}-\Pi^{s}_{|p_{y}|\to\infty}>0$, this implies that energy is transferred to the large-scale 2D flow, as depicted in Fig. 2.
At large-enough rotation, all modes are effectively in sector $H$. In this regime, all waves interacting with the large-scale 2D flow conserve their single-sign helicity: rotation has turned sign-definite an otherwise sign-indefinite invariant. This regime is obtained for $Ro_{\epsilon}\lesssim 0.1$ in Fig. 1(d), for which the transfer of energy is purely inverse, going from the waves to the large-scale 2D flow [26, 28]. In contrast, for $Ro_{\epsilon}\gtrsim 0.1$, heterochiral interactions are possible and generate a finite flux of energy to small scales, as we now explain.
Figure 2: (a) Energy transfer to the large-scale 2D flow in spectral space $(p_{y},p_{z})$ from QL theory and (b) schematic of the energy fluxes in the system. Orange line in (a) marks the boundary between Sectors H and A, (4). Waves in sector H conserve their sign-definite helicity $H_{\boldsymbol{p}}^{\pm}$, constraining the forward energy flux and causing energy transfer to the 2D flow. In contrast, modes in sector A break this conservation and extract energy from the 2D flow, with resulting energy flux $\Pi>\Pi^{*}$ (6), $\Pi^{*}$ being the energy input from sector H.
Energy flux in sector A — Modes in sector A are shear-dominated, and their energy balance is typical for a non-rotating flow. There, modes of opposite helicities interact, helicity is no longer conserved by sign, and energy is extracted from the shear flow. Using scale separation $l_{f}/L_{y}\ll 1$, the dynamics effectively reduces to that for a linear shear flow, which is analytically solvable when changing variables to fluctuating cross-flow velocity $v$ and vorticity $\eta$, see Eq. (12) and [20]. The condensate generates vorticity from $z$-varying $v$, which tilts the mean shear (via the RHS of Eq. (12b)), and converts mean-flow energy into turbulent kinetic energy (via Eq. (12c)).
Assuming an incoming flux of $\Pi^{*}(p_{x},p_{z})$ at $p=p^{*}$ into sector A, the flux $\Pi_{\boldsymbol{p}}\equiv\Pi^{+}_{\boldsymbol{p}}+\Pi^{-}_{\boldsymbol{p}}$ can be exactly computed (see End Matter), reading
$\Pi_{\boldsymbol{p}}$ $\underset{p_{y}\gg p^{*}_{y}}{=}\frac{\Pi^{*}}{2}\Bigg[1+\frac{p^{*2}}{p_{0}^{2}}\frac{p_{z}^{2}}{p_{x}^{2}}\Big(\frac{\pi}{2}-\arctan\Big(\frac{p_{y}^{*}}{p_{0}}\Big)\Big)^{2}\Bigg],$ (6)
@@
Total energy transfer and closure — To combine the results for sector H and A, we take the incoming flux $\Pi^{*}=\epsilon_{p_{x},p_{z}}(p_{x},p_{z})k_{f}/p^{*}$. The energy transfer to the 2D condensate as a function of the wavenumber $\boldsymbol{p}$ of inertial waves, ${U^{\prime}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$, can then be computed for a fixed value of ${U^{\prime}}/\Omega$ (see Eq. (19) and Fig. 2(a)). Waves in sector $H$ transfer energy to the condensate (${U^{\prime}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}>0$), producing a decaying energy flux, which flows into sector A, where waves extract energy from the condensate (${U^{\prime}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}<0$), see the schematic Fig. 2(b).
@@
Figure 3: (a) Energy transfer from 3D waves to 2D condensate and (b) rescaled amplitude of the condensate, as measured from the DNS and predicted from the QL theory. Dashed lines show the predictions in each regime: in black, Eq. (8); in red, the scalings derived in [26]. Solid lines show a numerical solution of the QL system at $Re=46$, including a wave cutoff $k_{\Omega}$ and a smooth treatment of homochiral-wave near-resonances. Inset in (b): rescaled amplitude of the 2D flow as a function of $Re$, showing the saturation to a flux-loop state.
@@
For $Ro_{\epsilon}\ll 1/b\sim 8l_{f}/L_{x}$ (here $\simeq 1.2$ when including subleading corrections), Eq.(8) gives $T_{\rm 3D-2D}\simeq\epsilon$ and $U^{\prime}\simeq\sqrt{\epsilon/\nu}$, corresponding to almost purely homochiral-wave interactions. This is consistent with the DNS behavior in Fig. 1 for $Ro_{\epsilon}{\color[rgb]{1,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{1,0,0}\sim}0.1$. For smaller $Ro_{\epsilon}$ the resonance condition (3) restricts these homochiral-wave interactions, leading to a progressive decoupling between 3D waves and the 2D flow, see the red dashed line in Fig. 3(a) obtained in [26]. The remaining energy is transferred to small-scale waves, which for $Ro_{\epsilon}\ll l_{f}/(2L_{y})$ is most of the injected energy. Note that the transition from the purely-homochiral to the mixed-interaction regime around $Ro_{\epsilon}\simeq 0.1$ in the theory (red and black dashed-lines in Fig. 3(a)) is smoothed out when using a smooth function of ${U^{\prime}}/\Omega$ (21) instead of the sharp near-resonant condition (3) for $\tilde{s}=s$. This perfects the agreement with the DNS in this region (solid line in Fig. 3(a)).
At the other end, for $Ro_{\epsilon}=\text{Ro}\times\text{Re}^{1/2}\gg 1/b$, Eq. (8) predicts that $U^{\prime}/\Omega\to 1/b$, a Reynolds-independent value. This asymptotic state corresponds to a vanishing energy transfer $T_{\rm 3D-2D}\to 0^{+}$. For fixed Ro and increasing Re in the DNS, ${U^{\prime}}/\Omega$ is indeed found to saturate to a constant value, closely following this prediction for the lowest $Ro$ (inset of Fig. 3b), and in accordance with previous observations of rotating turbulence under 3D [3] or 2D forcing [43].
The asymptotic flow state $T_{\rm 3D-2D}\underset{Re\to\infty}{\to}0^{+}$ with finite $U^{\prime}$ is dubbed a flux-loop state: a state where inverse and forward energy transfers are almost counterbalanced, see also [9, 19, 43, 15, 13]. This behavior, consistent with taking $\nu\to 0$ in the energy balance (1), would be impossible if the system strictly conserved two sign-definite quantities, as no steady state would be reached for $Re\to\infty$. Here, we reveal that the condensate self-tunes to a state of zero large-scale dissipation in this limit, achieved by an exact balance between the inverse energy transfer from modes conserving $H_{\boldsymbol{p}}^{\pm}$ (sector H), and the forward transfer due to modes breaking this conservation (sector A). Note that viscosity (and so Re) here primarily affects the large-scale dissipation, and could be replaced by, e.g., Ekman friction [33].
@@
When, instead, fixing Re and increasing Ro, the scaling $U^{\prime}\sim\Omega$ implies a vanishing 2D flow only for zero rotation $(Ro=\infty$), while condensates in DNS disappear to eddy turbulence at $Ro\simeq 0.5$, see Fig. 4. This is due to a shortage of waves at low rotation, and hence of modes in sector H, as we now explain.
When decreasing rotation, more modes behave as eddies instead of inertial waves, i.e. are such that $\tau_{nl}^{\rm eddy}<\tau_{\Omega}\equiv 2/\omega_{\boldsymbol{p}}$ (or, equivalently, $p>k_{\Omega}\equiv\Omega^{3/5}\epsilon^{-1/5}p_{z}^{3/5}$ [53]), with $\tau_{nl}^{\rm eddy}=\epsilon^{-1/3}p^{-2/3}$ the Kolmogorov eddy turnover time. Modes belonging to sector H are such that $\tau_{\Omega}\lesssim 1/{U^{\prime}}$, following Eq. (4). For $p>k_{\Omega}$, where they behave like eddies and $\tau_{nl}^{\rm eddy}<\tau_{\Omega}$, modes do not interact dominantly with the condensate as $\tau_{nl}^{\rm eddy}<1/{U^{\prime}}$, which breaks the QL approximation. As a consequence, the homochiral sector H, which is the source of energy for the condensate, is cutoff at $p=k_{\Omega}$ (sector A being cutoff at $k_{U}$ where eddy-eddy interactions dominate). As rotation is decreased, $k_{\Omega}$ diminishes and the energy input to the condensate is reduced, until no modes exist in sector H when $Ro=0.5$, for which $k_{\Omega}(p_{z}=k_{f})=k_{f}$. This marks the disappearance of condensates to a strong turbulence of 3D eddies. Introducing the wavenumber cutoff into our QL theory and integrating the resulting Reynolds stress numerically, we reproduce the drop in energy transfer and condensate amplitude observed in DNS, see the magenta solid lines in Fig. 3.
@@
While jets are found for most $Ro$ and $Re$, the condensate takes the form of cyclonic vortices at the boundary with 3D turbulence (stars in Figs. 1 and 4), previously observed in [36, 43, 13]. Our theory roughly captures the condensate amplitude for these states, but cannot explain scale selection nor symmetry breaking. The former may be due to time-scale competition between modes of different scales, consistent with the metastability of such states [18].
@@
We have untangled the organization of kinetic energy fluxes in rotating turbulence, showing it is controlled by helicity. Large-scale 2D structures are energized by fast-enough inertial waves via wave-dominated interactions which conserve helicity by sign. In contrast, the 2D flow loses energy to slow waves through shear-dominated interactions which break helicity-by-sign conservation, the waves carrying energy to small scales — similarly to non-rotating 3D turbulence. We have determined these bi-directional energy transfers analytically across rotation regimes via a quasi-linear wave-kinetic theory, which captures the $Ro$ and $Re$ behavior of 2D condensates in Navier-Stokes simulations, as summarized in Fig. 4.
Our results show that large-scale self-organization can be regulated by a sign-indefinite invariant in the presence of waves [7, 46], and thus arise even in systems that do not conserve multiple sign-definite quantities. The framework developed here provides a route to analyzing such behavior in other settings — such as thin layers [6], chiral flows [15], stratified flows [12, 44, 46, 32], magnetohydrodynamics [40, 22], plasmas [16] or geophysical flows [27] — and to elucidating the energy sinks and feedback mechanisms governing their self-organization.
@@ Acknowledgements.
We detail here the governing equations and solutions to the quasi-linear system describing the interactions between 3D inertial waves and a large-scale 2D mean flow. We consider a sinusoidal condensate of single mode $\boldsymbol{k}=\pm 2\pi/L_{y}$, $U_{\boldsymbol{k}}={U^{\prime}}/\sqrt{2}\delta_{\boldsymbol{k}-2\pi/L_{y}}$. The fluctuation field $\boldsymbol{u}(\boldsymbol{x})$ is decomposed into the helical basis $\boldsymbol{h}_{\boldsymbol{p}}^{s}=(\boldsymbol{p}\times[\boldsymbol{p}\times\mathbf{e}_{z}]-isp\boldsymbol{p}\times\mathbf{e}_{z})/(\sqrt{2}pp_{\perp})$, with $p=|\boldsymbol{p}|$ and $p_{\perp}=\sqrt{p_{x}^{2}+p_{y}^{2}}$. Following wave kinetic theory, we obtain evolution equations for the wave correlators $\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s^{\prime}}\rangle$, which we derive from the wave dynamical equation (2) in an expansion in ${U^{\prime}}/\Omega$ (see [26] and SM §B), and result in the helicity balance (5).
In Sector H, waves conserve their single-sign helicity, via Eq. (5) with $\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}^{s,-s}=0$. Introducing the correlator $\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}\equiv\frac{1}{2}(\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{s}\rangle)$, Eq. (5) is solved by the energy flux
$\Pi_{\boldsymbol{p}}^{s}\equiv\frac{-{U^{\prime}}p_{x}}{\sqrt{2}}\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s}=\left\{\begin{aligned} &~0~~\text{ if }\sigma p_{y}<-p_{y}^{f}\\ &\frac{\sigma\epsilon^{s}_{p_{x},p_{z}}}{2}\frac{k_{f}}{p}~~\text{ if }-p_{y}^{f}<\sigma p_{y}<p_{y}^{f}\\ &\sigma\epsilon^{s}_{p_{x},p_{z}}\frac{k_{f}}{p}~~\text{ if }\sigma p_{y}>p_{y}^{f},\end{aligned}\right$. (9)
@@
$\frac{{U^{\prime}}}{\sqrt{2}}\langle uv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ $=-\frac{{U^{\prime}}}{\sqrt{2}}\frac{p_{x}p_{y}}{p^{3}}(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{+}+\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-})$ (10)
which is strictly positive when $\sigma p_{y}>p_{y}^{f}$, and provided that the interaction is resonant at $p=k_{f}$: $\omega_{\boldsymbol{p}}^{s}+\omega_{-\boldsymbol{p}-\boldsymbol{k}}^{s}<{U^{\prime}}$.
In Sector A, the 2D flow mediates interactions between modes of opposite helicities, hence correlators $\Psi_{\boldsymbol{p}\boldsymbol{k}}^{s}\equiv\frac{1}{2}(\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}+\boldsymbol{k}}^{-s}\rangle+\langle a_{\boldsymbol{p}}^{s}a_{-\boldsymbol{p}-\boldsymbol{k}}^{-s}\rangle)$ are generated. They are determined from an evolution equation for $\langle a_{\boldsymbol{p}}^{s*}a_{\boldsymbol{p}}^{-s}\rangle$:
$\partial_{t}\langle a_{\boldsymbol{p}}^{s*}a_{\boldsymbol{p}}^{-s}\rangle=$ $\frac{{U^{\prime}}}{\sqrt{2}}\Bigg(\mathcal{L}_{\boldsymbol{p}}[\Psi_{\boldsymbol{p}\boldsymbol{k}}^{s,-s*}+\Psi_{\boldsymbol{p}\boldsymbol{k}}^{-s,s}])$
$~~~~~~-2V_{\boldsymbol{p}}^{s,-s*}(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s*}+\Phi_{\boldsymbol{p}\boldsymbol{k}}^{-s})\Bigg),$ (11)
where $\mathcal{L}_{\boldsymbol{p}}\equiv\Big[p_{x}\partial_{p_{y}}+\frac{p_{x}p_{y}}{p^{2}}-i\frac{p_{z}}{p}(1+2\frac{p_{x}^{2}}{p_{\perp}^{2}})\Big]$. Together with Eq. (5), Eq. (11) constitutes a closed system of correlators $(\Phi_{\boldsymbol{p}\boldsymbol{k}}^{s},\Psi_{\boldsymbol{p}\boldsymbol{k}}^{s})$. By changing variables from helical amplitudes $(a_{\boldsymbol{p}}^{+},a_{\boldsymbol{p}}^{-})$ to cross-flow velocity $v_{\boldsymbol{p}}$ and vorticity $\eta_{\boldsymbol{p}}\equiv(\nabla\times\boldsymbol{u})_{\boldsymbol{p},y}$, we obtain a triangular system of correlators from Eqs. (5)-(11), written in steady state as:
$\partial_{p_{y}}\Big(p^{4}\Re(\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}})\Big)=-p^{4}\frac{\sqrt{2}\epsilon\chi_{\boldsymbol{p},y}}{p_{x}{U^{\prime}}}$ (12a)
$\partial_{p_{y}}\Big(p^{2}\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}\Big)=ip^{2}\frac{p_{z}}{p_{x}}\langle vv\rangle_{\boldsymbol{p}\boldsymbol{k}}$ (12b)
$\partial_{p_{y}}\Re(\langle\eta\eta\rangle_{\boldsymbol{p}\boldsymbol{k}})=\frac{2p_{z}}{p_{x}}\Im\langle\eta v\rangle_{\boldsymbol{p}\boldsymbol{k}}-\frac{\sqrt{2}\epsilon p^{2}\chi_{\boldsymbol{p},y}}{p_{x}{U^{\prime}}},$ (12c)
with $\langle fg\rangle_{\boldsymbol{p}\boldsymbol{k}}\equiv(\langle f_{\boldsymbol{p}}g_{-\boldsymbol{p}+\boldsymbol{k}}\rangle+\langle f_{\boldsymbol{p}}g_{-\boldsymbol{p}-\boldsymbol{k}}\rangle)/2$. As the 3D modes have low frequencies in Sector A, rotation has no effect on the leading-order dynamics. Therefore, system (12) describes a non-rotating shear flow interacting with turbulent fluctuations.